OpenStudy (anonymous):

ok i really need some mor ehlpe please explain it Write the slope-intercept form of the equation of the line parallel to y = 5x + 2, which passes through the point (4, 9). please show the steps

hero (hero):

y = mx + b (9) = 5(4) + b 9 = 20 + b 9 - 20 = b -11 = b y = 5x - 11